Lim (3𝑥 − 𝑥 2 )


1
Expert's answer
2021-06-14T18:57:39-0400
"\lim\limits_{x\to-\infin}(3x-x^2)=\lim\limits_{x\to-\infin}(x)\lim\limits_{x\to-\infin}(3-x)=-\infin"

"\lim\limits_{x\to\infin}(3x-x^2)=\lim\limits_{x\to\infin}(x)\lim\limits_{x\to\infin}(3-x)=-\infin"

"\lim\limits_{x\to0}(3x-x^2)=3(0)-(0)^2=0"

"\lim\limits_{x\to a}(3x-x^2)=3a-a^2,\ a\not=\pm\infin"